# In-2-4D: Inbetweening from Two Single-View Images to 4D Generation

### v2

@@ Abstract
We pose a new problem, In-2-4D, for generative 4D (i.e., 3D + motion) inbetweening to interpolate two single-view images. In contrast to video/4D generation from only text or a single image, our interpolative task can leverage more precise motion control to better constrain the generation. Given two monocular RGB images representing the start and end states of an object in motion, our goal is to generate and reconstruct the motion in 4D, without making assumptions on the object category, motion type, length, or complexity. To handle such arbitrary and diverse motions, we utilize a foundational video interpolation model for motion prediction. However, large frame-to-frame motion gaps can lead to ambiguous interpretations. To this end, we employ a hierarchical approach to identify keyframes that are visually close to the input states while exhibiting significant motions, then generate smooth fragments between them. For each fragment, we construct a 3D representation of the keyframe using Gaussian Splatting (3DGS). The temporal frames within the fragment guide the motion, enabling their transformation into dynamic 3DGS through a deformation field. To improve temporal consistency and refine the 3D motion, we expand the self-attention of multi-view diffusion across timesteps and apply rigid transformation regularization. Finally, we merge the independently generated 3D motion segments by interpolating boundary deformation fields and optimizing them to align with the guiding video, ensuring smooth and flicker-free transitions. Through extensive qualitative and quantitive experiments as well as a user study, we demonstrate the effectiveness of our method and design choices. The code will be made public upon acceptance
1 Introduction
@@
Figure 1: In-2-4D: 4D motion inbetweening from 2 single-view images, a minimalistic input setup while still offering visual motion control. Given two monocular RGB images of an object at two distinct motion states (start and end), our method generates a smooth, natural, and seamless 4D (3D object + motion) interpolation between them. We make no assumptions on the object category or motion type. Piecewise rigid, freeform, and long-range motions, even with topology changes (e.g., egg dropping in liquid causing a splash), are all allowed.
@@
Motion inbetweening is a classic animation problem. When generating motions of 3D objects, the typical input consists of a 3D object in two distinct motion states, e.g., as in point cloud interpolation [27, 49]. With significant progresses in 3D generative AI in recent years, many recent attempts have been made on “video-to-4D” [48, 41, 30, 47, 11], whose task is to “lift” an object captured in a video into the 3D space so its motion from the video can be viewed from all angles. However, videos can be unrealistic (e.g., for very slow motions such as the flower blossoming in Fig. 1) or at least expensive (e.g., for very fast motions) to acquire. In other application settings such as motion planning and visual storyboarding [50], the motions produced are the outcomes to be explored; they were not acquirable in the first place.
@@
An intriguing question is whether the inbetweening and lifting problems above can be “fused” to produce 4D contents, i.e., 3D object with motion, from a minimalistic and easy-to-acquire input setup. Indeed, with rapid advances in video foundation models and a push for controllable motion generation, there has been an emergence of works on video frame interpolation [29, 8, 6, 44, 38] which produce video, rather than 4D, outputs from two images.
@@
In this paper, we seek a solution to the novel task of generating 4D interpolative contents from merely two single-view images capturing an object in two distinct motion states. Compared to video or 4D generation from text only [35, 1] or from a single image [30], our interpolative task anchored on two images can leverage more precise motion control to better constrain the generative process and lead to more plausible and cleaner results, as shown in Figs. 1 and 2.
@@
Figure 2: Comparing to text-to-4D and single-image-to-4D. Text-only and single-image inputs are insufficient to constrain the 4D generation by state-of-the-art models (L4GM [31, 48] , leading to unnatural results and strong artifacts, especially for complex motions such as the one shown. Results from our method can be found in Fig. 1.
@@
Figure 3: Illustration of our In-2-4D solution pipeline. Given two single-view images as input (a), we first generate keyframes inbetween them to bridge the motion gaps so as to avoid abrupt motions between consecutive keyframes to generate video fragments (b) . These keyframes are then utilized to learn static 3D geometries per fragment which are then deformed using a deformation field (e.g., Hexplane) to obtain a 4D scene in the form of 3D Gaussian splats (3DGS) per fragment. Finally, we obtain a 4D scene in the form of 3D Gaussian splats (3DGS) per fragment. (c) To aggregate the deformations, we linearly interpolate the deformation fields (d) in a cascading fashion and apply smoothing constraints on the 3DGS to improve geometric and structural consistency between the generated novel views.
@@
At the high level, our method operates in two phases: 2D still images to video via interpolation, and then video-to-4D via lifting, as illustrated in Fig. 3. To handle arbitrary and diverse motions, we leverage video foundational models. However, most such models are built on video diffusion [3], which has been trained predominantly by short videos. As such, they can be ineffective for motion inbetweening when the input states span large geometry or structural changes, resulting in large motion “jumps” and absence of detailed and intricate object movements. When such video models are used to model complex motion, they often suffer from visible artifacts which further detoriate the novel views when lifted to 4D space.
@@
To this end, we develop a divide-and-conquer approach to adaptively and recursively generate a set of keyframes to form a coarse outline of the motion, breaking the long-range and nonlinear motion into simpler, temporally-local transitions that are easier to model and interpolate. To start, we employ a foundational video interpolation model such as DynamiCrafter [43] to generate an initial set of intermediate frames between the two input states. Then we perform motion and appearance feature analyses over these frames to select one or more keyframes that are visually close to the input states and show significant motion jumps. Consecutive keyframes that incur a large motion will anchor a new video interpolation to generate more immediate frames, effectively “magnifying” the said motion. This process is carried out hierarchically until all motions between consecutive keyframes are sufficiently small. Notably, our video fragment generation is training-free and requires no fine-tuning of video diffusion models, making it both accessible and effective for 4D generation. For each video fragment, and in parallel with other fragments, we first learn a distinct static 3D Gaussian splatting (3DGS) model to capture the object geometry. We then apply a deformation field to convert this 3DGS into a dynamic, i.e., 4D, model by utilizing multi-view diffusion priors to refine the warping, geometry, and textures over unseen areas. By construction, the fragment contains relatively simple motions, hence multi-view generation can effectively mitigate texture degradation and geometry misalignment.
@@
We conduct extensive experiments on I4D-15 for evaluation. Quantitative and qualitative comparisons are made to methods and baselines to demonstrate the effectiveness of our method in terms of the quality of generated results, generalizability, and handling of a variety of motions; see Fig. 1. While achieving superior generation quality and remaining faithful to the input images than all current alternatives, our solution is far from artifact-free. This highlights the significant challenges that still lie ahead with 4D content generation from highly sparse inputs which still aim to exert intricate control. As a first attempt at tackling such a complex problem, we hope it can serve as a promising start to stimulate future work.
2 Related Work
Video inbetweening. Recent methods have extended pre-trained diffusion-based text-to-image models to generate motion from static images by adapting UNets to temporal data [40, 15, 34]. One notable model is AnimateDiff [9], which learns low-rank adapters for diverse motion patterns. More recent approaches condition pre-trained text-to-video models on input images. VideoCrafter1 [5] uses dual cross-attention layers to combine image features with text prompts, while DynamiCrafter [43] further refines this by concatenating the input image with noisy latent features. Our method builds on DynamiCrafter to enhance its outputs through recursive video magnification. While several video magnification techniques exist [16, 23], we leverage a video inbetweening network (e.g., DynamiCrafter) to interpolate frames and amplify motion when large displacements are present. Decomposing large motions into smaller fragments with smoother transitions reduces geometric ambiguities between consecutive frames, producing 4D results with fewer artifacts and improved visual quality.
4D scene interpolation. Dynamic 3D scene interpolation (4D Interpolation) is recently becoming more popular in 4D literature. Earlier works [26] leverage neural radiance fields (NeRF) for temporally coherent 3D reconstructions, while NeuralPCI [49] employs neural fields for multi-frame, non-linear 3D point cloud interpolation. PAPR [27] estimates motion via point-based rendering and local displacement optimization. Recent methods [22, 11, 31] use frame motions from Diffusion-based Video Interpolation models [2, 43] to infer 3D deformation. However, Video Diffusion models [3], trained on short clips (e.g., 16 frames), struggle with long sequences, causing artifacts from large per-frame motion jumps. To address this, we use generative Video Interpolation models (e.g., DynamiCrafter [43]) hierarchically for longer 3D trajectory estimation without extra training.
@@ 2 Related Work
4D dynamic scene generation. Recent works [39, 46, 20, 10, 48, 30, 41] extend 3D Gaussian Splatting (3DGS)[13] to 4D using time-conditioned deformation networks with SDS and multi-view geometry. MAV3D [35] pioneered text-to-4D via NeRF and score distillation, followed by similar approaches [1]. Consistent4D [11] introduces video-to-4D with pre-trained image diffusion models, extending to image/video-conditional 4D generation [33, 37]. STAG4D [48] and 4DGen [47] refine diffusion with pseudo-labels, while SC4D [41] employs sparse Gaussians and LBS for dynamic 3D. L4GM [31] proposed a 4D foundation model effective for simple motions. Despite progress, video-to-4D methods struggle with high dynamics, accumulating errors over long videos due to reliance on a single canonical model. We mitigate this by segmenting videos into shorter fragments with their own canonical model to improving geometric consistency.
3 Methodology
@@
An overview of our method is shown in Fig. 3. Given two images representing the start and end states of an object in motion, we aim to generate and reconstruct the motion in 4D (3D+motion). To predict the motion, we use a video interpolation model, but large motions between frames can lead to ambiguous interpretations and results with artifacts. To solve this, we employ a hierarchical approach to identify keyframes that are visually close to the input states and exhibit significant motion, then generate smooth fragments between them. For each fragment, the 3D representation of the keyframe is first constructed using Gaussian Splatting. The temporal frames within the fragment serve as motion guidance, enabling their transformation into dynamic Gaussians through a deformation field. To enhance temporal consistency and refine 3D motion of the fragment, we expand the self-attention of multi-view diffusion across time steps and introduce rigid transformation regularization. Finally, the independently generated 3D motion segments are merged by interpolating the boundary deformation fields and optimizing them to align with the guiding video. This ensures smooth and flicker-free transitions.
3.1 Problem Setup
@@
Figure 4: Illustration of Hierarchal Fragment Generation. At each generation step, a keyframe is selected by finding the largest motion from the DIFT heatmap and FID score. New frames are re-generated using the keyframe to minimize large motion changes. Selection of the keyframes and re-generation is done in a hierarchial manner to generate fragments having simple motions
3.2 Temporal Fragment Hierarchy
We propose a method for identifying keyframes in fragments with significant deformations and adaptively expanding them. Large deformations between start and end states induce rapid intermediate motion changes, which hinder 3D deformation learning [21] as shown in Fig. 5. To mitigate this, we partition the motion trajectory into fragments with smoother quasi-static motions, selecting keyframes densely in dynamic regions and sparsely in static regions. This balances training overhead, model size, and performance and enhances temporal consistency.
Hierarchical key-frame generator. To generate keyframes for the intermediate motion between two initial states, we employ a Video Interpolation Model (e.g., DynamiCrafter), denoted as $\psi(.)$. Given input images $I_{s}$ and $I_{e}$ along with a motion prompt $p$ (extracted using BLIP [17]), we generate a sequence of latent frames $Z$. The pairwise DIFT [36] features quantify frame-wise similarity, enabling a rapid assessment of motion changes. As illustrated in Fig. 4, a heatmap visualizes temporal variations, where significant object movements or new appearances are represented as bright regions. The heatmap between frames $I_{i}$ and $I_{j}$ is computed as:

### v3

@@ Abstract.
We pose a new problem, In-2-4D, for generative 4D (i.e., 3D + motion) inbetweening to interpolate two single-view images. In contrast to video/4D generation from only text or a single image, our interpolative task can leverage more precise motion control to better constrain the generation. Given two monocular RGB images representing the start and end states of an object in motion, our goal is to generate and reconstruct the motion in 4D, without making assumptions on the object category, motion type, length, or complexity. To handle such arbitrary and diverse motions, we utilize a foundational video interpolation model for motion prediction. However, large frame-to-frame motion gaps can lead to ambiguous interpretations. To this end, we employ a hierarchical approach to identify keyframes that are visually close to the input states while exhibiting significant motions, then generate smooth fragments between them. For each fragment, we construct a 3D representation of the keyframe using Gaussian Splatting (3DGS). The temporal frames within the fragment guide the motion, enabling their transformation into dynamic 3DGS through a deformation field. To improve temporal consistency and refine the 3D motion, we expand the self-attention of multi-view diffusion across timesteps and apply rigid transformation regularization. Finally, we merge the independently generated 3D motion segments by interpolating boundary deformation fields and optimizing them to align with the guiding video, ensuring smooth and flicker-free transitions. Through extensive qualitative and quantitive experiments as well as a user study, we demonstrate the effectiveness of our method and design choices.
Project Page & Source Code: https://in-2-4d.github.io/
3D motion interpolation, diffusion models, 4D reconstruction and synthesis from images
submissionid: 1660journal: TOGjournalyear: 2025copyright: ccconference: SIGGRAPH Asia 2025 Conference Papers; December 15–18, 2025; Hong Kong, Hong Kongbooktitle: SIGGRAPH Asia 2025 Conference Papers (SA Conference Papers ’25), December 15–18, 2025, Hong Kong, Hong Kongdoi: 10.1145/3757377.3763904isbn: 979-8-4007-2137-3/2025/12
1. Introduction
@@
Figure 1. In-2-4D: 4D motion inbetweening from 2 single-view images, a minimalistic input setup while still offering visual motion control. Given two monocular RGB images of an object at two distinct motion states (start and end), our method generates a smooth, natural, and seamless 4D (3D object + motion) interpolation between them. We make no assumptions on the object category or motion type. Piecewise rigid, freeform, and long-range motions, even with topology changes (e.g., egg dropping in liquid causing a splash), are all allowed.
@@
Motion inbetweening is a classic animation problem. When generating motions of 3D objects, the typical input consists of a 3D object in two distinct motion states, e.g., as in point cloud interpolation (Peng et al., 2024; Zheng et al., 2023). With significant progresses in 3D generative AI in recent years, many recent attempts have been made on “video-to-4D” (Zeng et al., 2024; Wu et al., 2024; Ren et al., 2023; Yin et al., 2023; Jiang et al., 2023), whose task is to “lift” an object captured in a video into the 3D space so its motion from the video can be viewed from all angles. However, videos can be unrealistic (e.g., for very slow motions such as the flower blossoming in Fig. 1) or at least expensive (e.g., for very fast motions) to acquire. In other application settings such as motion planning and visual storyboarding (Zhong et al., 2025), the motions produced are the outcomes to be explored; they were not acquirable in the first place.
@@
An intriguing question is whether the inbetweening and lifting problems above can be “fused” to produce 4D contents, i.e., 3D object with motion, from a minimalistic and easy-to-acquire input setup. Indeed, with rapid advances in video foundation models and a push for controllable motion generation, there has been an emergence of works on video frame interpolation (Reda et al., 2022; Feng et al., 2024; Danier et al., 2024; Xing et al., 2024; Wang et al., 2025) which produce video, rather than 4D, outputs from two images.
@@
In this paper, we seek a solution to the novel task of generating 4D interpolative contents from merely two single-view images capturing an object in two distinct motion states. Compared to video or 4D generation from text only (Singer et al., 2023; Bahmani et al., 2024) or from a single image (Ren et al., 2023), our interpolative task anchored on two images can leverage more precise motion control to better constrain the generative process and lead to more plausible and cleaner results, as shown in Figs. 1 and 2.
@@
Figure 2. Comparing to text-to-4D and single-image-to-4D. Text-only and single-image inputs are insufficient to constrain the 4D generation by state-of-the-art models (L4GM (Ren et al., 2025; Zeng et al., 2024) , leading to unnatural results and strong artifacts, especially for complex motions such as the one shown. Results from our method can be found in Fig. 1.
@@
Figure 3. Illustration of our In-2-4D solution pipeline. Given two single-view images as input (a), we first generate keyframes inbetween them to bridge the motion gaps so as to avoid abrupt motions between consecutive keyframes to generate video fragments (b) . These keyframes are then utilized to learn static 3D geometries per fragment which are then deformed using a deformation field (e.g., Hexplane) to obtain a 4D scene in the form of 3D Gaussian splats (3DGS) per fragment. Finally, we obtain a 4D scene in the form of 3D Gaussian splats (3DGS) per fragment. (c) To aggregate the deformations, we linearly interpolate the deformation fields (d) in a cascading fashion and apply smoothing constraints on the 3DGS to improve geometric and structural consistency between the generated novel views.
@@
At the high level, our method operates in two phases: 2D still images to video via interpolation, and then video-to-4D via lifting, as illustrated in Fig. 3. To handle arbitrary and diverse motions, we leverage video foundational models. However, most such models are built on video diffusion (Blattmann et al., 2023b), which has been trained predominantly by short videos. As such, they can be ineffective for motion inbetweening when the input states span large geometry or structural changes, resulting in large motion “jumps” and absence of detailed and intricate object movements. When such video models are used to model complex motion, they often suffer from visible artifacts which further detoriate the novel views when lifted to 4D space.
@@
To this end, we develop a divide-and-conquer approach to adaptively and recursively generate a set of keyframes to form a coarse outline of the motion, breaking the long-range and nonlinear motion into simpler, temporally-local transitions that are easier to model and interpolate. To start, we employ a foundational video interpolation model such as DynamiCrafter (Xing et al., 2023) to generate an initial set of intermediate frames between the two input states. Then we perform motion and appearance feature analyses over these frames to select one or more keyframes that are visually close to the input states and show significant motion jumps. Consecutive keyframes that incur a large motion will anchor a new video interpolation to generate more immediate frames, effectively “magnifying” the said motion. This process is carried out hierarchically until all motions between consecutive keyframes are sufficiently small. Notably, our video fragment generation is training-free and requires no fine-tuning of video diffusion models, making it both accessible and effective for 4D generation. For each video fragment, and in parallel with other fragments, we first learn a distinct static 3D Gaussian splatting (3DGS) model to capture the object geometry. We then apply a deformation field to convert this 3DGS into a dynamic, i.e., 4D, model by utilizing multi-view diffusion priors to refine the warping, geometry, and textures over unseen areas. By construction, the fragment contains relatively simple motions, hence multi-view generation can effectively mitigate texture and geometry degradation.
@@
We conduct extensive experiments on I4D-15 and Consistent4D (Jiang et al., 2023) benchmark for evaluation. Quantitative and qualitative comparisons are made to methods and baselines to demonstrate the effectiveness of our method in terms of the quality of generated results, generalizability, and handling of a variety of motions; see Fig. 1. While achieving superior generation quality and remaining faithful to the input images than all current alternatives, our solution is far from artifact-free. This highlights the significant challenges that still lie ahead with 4D content generation from highly sparse inputs which still aim to exert intricate control. As a first attempt at tackling such a complex problem, we hope it can serve as a promising start to stimulate future work.
2. Related Work
In computer graphics, most works on motion inbetweening have been developed for character animation, e.g., (Zhou et al., 2020; Cohan et al., 2024; Qin et al., 2022; Zhong et al., 2025). Our work aims to handle arbitrary and diverse motions, leveraging foundational models, for the novel task of In-2-4D. In this section, we mainly cover related works on 4D generation and video/4D interpolation.
@@ 2. Related Work
4D dynamic scene generation. Recent works (Wu et al., 2023b; Yang et al., 2023; Liang et al., 2023; Guo et al., 2024a; Zeng et al., 2024; Ren et al., 2023; Wu et al., 2024) extend 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023a) to 4D using time-conditioned deformation networks with Score Distillation Sampling (SDS) and multi-view geometry. MAV3D (Singer et al., 2023) pioneered text-to-4D via NeRF and score distillation, followed by similar approaches (Bahmani et al., 2024). Consistent4D (Jiang et al., 2023) introduces video-to-4D with pre-trained image diffusion models, extending to video-conditional 4D generation (Shi et al., 2023; Wang and Shi, 2023; Yang et al., 2024b). STAG4D (Zeng et al., 2024) and 4DGen (Yin et al., 2023) refine diffusion with pseudo-labels, while SC4D (Wu et al., 2024) employs sparse Gaussians and Linear Binding Skinning (LBS) for dynamic 3D generation. Recently, research has shifted towards SDS-free 4D foundation models (Ren et al., 2025; Xie et al., 2024; Wu et al., 2025; Zhang et al., 2024; Jiang et al., 2024; Liang et al., 2024b; Sun et al., 2024) that synthesize synchronized multi-view videos as intermediates for 4D reconstruction, bypassing SDS optimization. While these models produce plausible 4D motions, they remain limited by the inductive bias of their training data and fail to generalize to scenes with complex dynamics. Despite progress in video diffusion, current video-to-4D methods still struggle with highly dynamic motions and accumulate errors over long sequences due to reliance on a single canonical model. To address this, our framework segments videos into shorter fragments, via keyframe generation, each associated with its own 3D canonical model, thereby improving geometry.
Video inbetweening or frame interpolation. Motion inbetweening allows finer-grained control compared to textual inputs or motion extrapolation from a single frame. Recent methods have extended pre-trained diffusion-based text-to-image models to generate motion from static images by adapting UNets to temporal data (Wu et al., 2023a; Khachatryan et al., 2023; Singer et al., 2022). One notable model is AnimateDiff (Guo et al., 2024b), which learns low-rank adapters for diverse motion patterns. More recent approaches condition pre-trained text-to-video models on input images. VideoCrafter1 (Chen et al., 2023) uses dual cross-attention layers to combine image features with text prompts, while DynamiCrafter (Xing et al., 2023) further refines this by concatenating the input image with noisy latent features. Our method builds on DynamiCrafter to enhance its outputs through recursive video magnification. While several video magnification techniques exist (Le Ngo and Phan, 2019; Liu et al., 2005), we leverage a video inbetweening network (e.g., DynamiCrafter) to interpolate frames and amplify motion when large displacements are present. Decomposing large motions into smaller fragments with smoother transitions reduces geometric ambiguities between consecutive frames, producing 4D results with fewer artifacts and improved visual quality.
4D scene interpolation. Lately, dynamic 3D scene interpolation, or 4D Interpolation, has become more popular. Earlier works (Park et al., 2023) leverage neural radiance fields (NeRF) for temporally coherent 3D reconstructions, while NeuralPCI (Zheng et al., 2023) employs neural fields for multi-frame, non-linear 3D point cloud interpolation. PAPR (Peng et al., 2024) estimates motion via point-based rendering and local displacement optimization. Recent methods (Ling et al., 2024; Jiang et al., 2023; Ren et al., 2025) use frame motions from Diffusion-based Video Interpolation models (Blattmann et al., 2023a; Xing et al., 2023) to infer 3D deformation. However, Video Diffusion models (Blattmann et al., 2023b), trained on short clips (e.g., 16 frames), struggle with long sequences, causing artifacts from large per-frame motion jumps. To address this, we use generative Video Interpolation models (e.g., DynamiCrafter (Xing et al., 2023)) hierarchically for longer 3D trajectory estimation without extra training. In addition, In-2-4D represents the first attempt at general 4D interpolation from two single-view motion frames.
3. Methodology
@@
An overview of our method is shown in Fig. 3. Given two images representing the start and end states of an object in motion, we aim to generate and reconstruct the motion in 4D (3D+motion). To predict the motion, we use a video interpolation model, but large motions between frames can lead to ambiguous interpretations and results with artifacts. To solve this, we employ a hierarchical approach to identify keyframes that are visually close to the input states and exhibit significant motion, then generate smooth fragments between them. For each fragment, the 3D representation of the keyframe is first constructed using Gaussian Splatting. The temporal frames within the fragment serve as motion guidance, enabling their transformation into dynamic Gaussians through a deformation field. To enhance temporal consistency and refine 3D motion of the fragment, we expand the self-attention of multi-view diffusion across time steps and introduce rigid transformation regularization. Finally, the independently generated 3D motion segments are merged by interpolating the boundary deformation fields and optimizing them to align with the guiding video. This ensures flicker-free transitions.
3.1. Problem Setup
@@
Figure 4. Illustration of Hierarchal Fragment Generation. At each generation step, a keyframe is selected by finding the largest motion from the DIFT heatmap and FID score. New frames are re-generated using the keyframe to minimize large motion changes. Selection of the keyframes are done in a hierarchial manner to generate fragments having simple motions
3.2. Temporal Fragment Hierarchy
We propose a method for identifying keyframes in fragments with significant deformations and adaptively expanding them. Large deformations between start and end states induce rapid intermediate motion changes, which hinder 3D deformation learning (Liang et al., 2024a) as shown in Fig. 5. To mitigate this, we partition the motion trajectory into fragments with smoother quasi-static motions, selecting keyframes densely in dynamic regions and sparsely in static regions. This balances training overhead, model size, and performance and enhances temporal consistency.
Hierarchical key-frame generator. To generate keyframes for the intermediate motion between two initial states, we employ a Video Interpolation Model (e.g., DynamiCrafter), denoted as $\psi(.)$. Given input images $I_{s}$ and $I_{e}$ along with a motion prompt $p$ (extracted using BLIP (Li et al., 2022)), we generate a sequence of latent frames $Z$. The pairwise DIFT (Tang et al., 2023) features quantify frame-wise similarity, enabling a rapid assessment of motion changes. As illustrated in Fig. 4, a heatmap visualizes temporal variations, where significant object movements or new appearances are represented as bright regions. The heatmap between frames $I_{i}$ and $I_{j}$ is computed as: